# Neural Audio Codecs for Prompt-Driven Universal Source Separation

### v2

@@ 5 Conclusion
Limitations and clarifications. (1) Codec/bandwidth. Our experiments use a 16 kHz mono NAC for reproducibility and on-device feasibility; however, CodecSep is codec-agnostic. The masker operates on continuous NAC latents with a frozen codec, so upgrading to 44.1/48 kHz and/or stereo is a drop-in backbone swap with no architectural changes to masking or conditioning. (2) Data and prompts. Training data scale and prompt diversity are modest relative to open-domain audio. As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains. (3) Temporal prompting. While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction. (4) Perceptual SFX quality. In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work.
Future Work. We will (i) scale the frozen codec backbone to higher bandwidths and stereo (44.1/48 kHz), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
@@ Appendix C Training Details
Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep. For AudioSep, we evaluate the publicly available pre-trained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps using our setup for consistency. We also include SDCodec, using the official pre-trained models made available by the authors. To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing during training and inference. This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
@@
Table 7: Generalization and transfer results for universal source separation.
@@
Appendix F Further Studies: Reconstruction Performance.
@@
Table 9: Results: Reconstruction Performance, Universal Source Separation (dnr-v2-test)

### v3

@@ 5 Conclusion
Limitations and clarifications. (1) Data and prompts. Training data scale and prompt diversity are modest relative to open-domain audio. As shown in Table 2, finer SFX supervision sharpens SFX extraction and improves speech/music stems; larger, more heterogeneous corpora spanning multiple prompt granularities—including temporal/relational cues—should yield further gains. (2) Temporal prompting. While CodecSep is robust to synonymic paraphrases, we did not evaluate prompts with explicit temporal structure (e.g., causal ordering), which remains an open direction. (3) Perceptual SFX quality. In some settings, SFX perceptual quality trails the best competing scores despite superior SI-SDR; improving SFX naturalness without sacrificing separation is future work.
Future Work. We will (i) extend separation to higher bandwidths and stereo/spatial audio (e.g., 48 kHz stereo Encodec; HO–DirAC Hold et al. (2024), SpatialCodec Xu et al. (2024)), (ii) broaden prompt coverage with temporal/relational and referring-expression cues, (iii) explore stronger separator objectives, (iv) extend conditioning beyond text to audio/image and mixed prompts via the same FiLM interface.
@@ Appendix C Training Details
Since TDANet and CodecFormer were originally developed for speech separation, we re-trained newly initialized versions on the dnr-v2 training set using the same configuration as CodecSep. For AudioSep, we evaluate the publicly available pre-trained model—trained on diverse datasets—and versions re-trained on dnr-v2 and AudioCaps using our setup for consistency. We also include SDCodec, using the official pre-trained models made available by the authors. To ensure a fair comparison, all inputs to TDANet and AudioSep undergo codec processing with a full-band stereo-capable $48$ kHz EnCodec during training and inference. This accounts for codec-induced distortions and artifacts, reflecting realistic deployment scenarios where audio is typically processed through compression pipelines in cloud-based systems.
@@
Table 7: Generalization and transfer results for universal sound separation.
@@
Appendix F Bandwidth scaling: extending CodecSep to 48 kHz full-band
[…]
Table 9: Results: Extending CodecSep to $48$ kHz full-band (dnr-v2-test)
Model Sampling Rate Metric ($\uparrow$) Music Speech Sfx
AudioSep $32$ kHz SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
AudioSep + dnr-v2 $32$ kHz SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
CodecSep + dnr-v2 $16$ kHz SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
(DAC Backbone) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
CodecSep + dnr-v2 $48$ kHz SI-SDR $2.8^{\pm 3.5}$ $5.36^{\pm 2.36}$ $0.46^{\pm 3.83}$
(EnCodec Backbone) ViSQOL $2.39^{\pm 0.5}$ $2.63^{\pm 0.41}$ $2.37^{\pm 0.65}$
Appendix G Further Studies: Reconstruction Performance.
@@ Appendix F Bandwidth scaling: extending CodecSep to 48 kHz full-band
Table 10: Results: Reconstruction Performance, Universal Sound Separation (dnr-v2-test)